### v1

# Highly Anisotropic Charge Dynamics and Spectral Weight Redistribution in the Trilayer Nickelate La4Ni3O10

Abstract
We study the abab-plane and $c$-axis charge dynamics of La4Ni3O10 using optical spectroscopy. While a pronounced Drude profile, i.e. metallic response, is observed in the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$, the $c$-axis optical spectra $\sigma_{1}^{c}(\omega)$ exhibit semiconducting behavior. The zero-frequency extrapolation of the optical conductivity $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$ gives a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K for La4Ni3O10, which is much larger than the values in iron-based superconductors but comparable to those in high-$T_{c}$ cuprates. The interband response is also highly anisotropic, showing salient orbital selectivity for light polarized in the abab plane and along the $c$ axis. The interband-transition peaks in both $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ are located at lower energies compared to density-functional-theory predictions, signifying considerable electronic correlations. By investigating the spectral weight transfer, we find that in the pristine phase, Coulomb correlations have a marked impact on the charge dynamics of La4Ni3O10, whereas in the density-wave state, a gap opens with the Ni-$d_{z^{2}}$ orbital being involved.
[…]
Figure 1: (a) Reflectivity of La4Ni3O10 at different temperatures for the abab plane $R^{ab}(\omega)$ and $c$ axis $R^{c}(\omega)$. The inset shows the temperature dependence of $R^{ab}(\omega)$ (red solid circle) and $R^{c}(\omega)$ (blue solid diamond) at $\omega$ = 1000 $\textrm{cm}^{-1}$. (b) The abab-plane optical conductivity $\sigma^{ab}_{1}(\omega)$ and $c$-axis optical conductivity $\sigma^{c}_{1}(\omega)$ of La4Ni3O10 at different temperatures. The inset displays $\sigma^{ab}_{1}(\omega)$ and $\sigma^{c}_{1}(\omega)$ at 300 K on the logarithmic scale with the zero-frequency extrapolations being denoted by solid circles. (c) and (d) show $R(\omega)$ and $\sigma_{1}(\omega)$ at 300 K along different directions in a broad frequency range up to 25 000 $\textrm{cm}^{-1}$, respectively.
We investigate the abab-plane and $c$-axis optical properties of La4Ni3O10. In the abab plane, we found metallic behavior, as evidenced by a conspicuous Drude profile in the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$. In sharp contrast, the absence of a Drude response in the $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ indicates semiconducting behavior along the $c$ axis. Since $\sigma_{1}(\omega\rightarrow 0)\equiv 1/\rho_{\text{dc}}$, by extrapolating $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ to zero frequency, we obtained $\rho_{c}/\rho_{ab}\simeq 366$ at 300 K. Such a large resistivity anisotropy is reminiscent of high-$T_{c}$ cuprates. Light polarized in the abab plane and along the $c$ axis selectively excites interband transitions between different orbitals, giving rise to a strong anisotropy of the interband response. A detailed examination of the spectral weight redistribution in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ reveals that the charge dynamics in the pristine phase is strongly influenced by Coulomb correlations, whereas the density-wave state is characterized by the opening of a gap which involves the Ni-$d_{z^{2}}$ orbital.
Sample growth, characterization, optical measurements and data analysis methods are detailed in the Supplemental Materials Sup . Transport measurements indicate that the intertwined charge- and spin-density-wave transition occurs at $T_{\text{DW}}$ = 136.5 K in our La4Ni3O10 sample (Supplemental Materials Sup ). Figure 1(a) shows the reflectivity of La4Ni3O10 for light polarized in the abab plane ($E\parallel ab$) $R^{ab}(\omega)$ and along the $c$ axis ($E\parallel c$) $R^{c}(\omega)$ up to 2000 $\textrm{cm}^{-1}$ at several representative temperatures. $R^{ab}(\omega)$ is high and approaches unity in the low-frequency limit. Additionally, it increases as $T$ is lowered. All these properties suggest a metallic nature of La4Ni3O10 in the abab plane. Below $T_{\text{DW}}$, $R^{ab}(\omega)$ is significantly suppressed in a broad frequency range above 500 $\textrm{cm}^{-1}$ due to the density-wave gap opening, in agreement with a previous optical study Xu et al. (2025). By contrast, $R^{c}(\omega)$ attains a constant value much lower than 1 ($\sim$0.6) in the low-frequency limit and is dominated by a series of sharp features arising from IR-active phonon modes in the far-infrared range, which are prototypical characteristics of semiconducting materials Jiang et al. (2020); Hao et al. (2021); Lobo et al. (2007). The inset of Fig. 1(a) plots the $T$ dependence of $R^{ab}(\omega)$ (red solid circles) and $R^{c}(\omega)$ (blue solid diamonds) at $\omega$ = 1000 $\textrm{cm}^{-1}$. A drop is observed below $T_{\text{DW}}$ in both $R^{ab}(\omega)$ and $R^{c}(\omega)$, indicating that both the abab-plane and $c$-axis charge dynamics are affected by the density-wave transition in La4Ni3O10. Figure 1(c) displays $R^{ab}(\omega)$ (red solid curve) and $R^{c}(\omega)$ (blue solid curve) at 300 K up to 25 000 $\textrm{cm}^{-1}$, highlighting the strong anisotropy of La4Ni3O10 in a broad frequency range.
Figure 2: (a) The measured $\sigma_{1}^{ab}(\omega)$ at 150 K (cyan solid curve) and the Drude-Lorentz fit (black dashed line). The fit is decomposed into two Drude components D1 (red shaded area), D2 (blue shaded area), and a series of Lorentz components L1ab (green hatched area), L2ab (orange hatched area), L3ab (violet hatched area) and LHab (grey hatched area). (b) The measured $\sigma_{1}^{c}(\omega)$ at 150 K (cyan solid line) and the fit (black dashed line), which consists of several Lorentz components L1c (green hatched area), L2c (orange hatched area) and LHc (grey hatched area). (c) and (d) show the calculated interband $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$, respectively. The vertical dashed lines mark the positions of interband-transition peaks. (e) The calculated electronic band structure of La4Ni3O10, with magenta and navy denoting the Ni-$d_{z^{2}}$ and Ni-$d_{x^{2}-y^{2}}$ orbitals, respectively. Energy- and $\vec{k}$-resolved transition dipole moments for (f) $E\parallel ab$ and (g) $E\parallel c$. The initial and final states are colored with red and blue, respectively.
Figure 1(b) depicts the abab-plane optical conductivity $\sigma_{1}^{ab}(\omega)$ and $c$-axis optical conductivity $\sigma_{1}^{c}(\omega)$ of La4Ni3O10 up to 2000 $\textrm{cm}^{-1}$ at various temperatures. While the low-frequency $\sigma_{1}^{ab}(\omega)$ is characterized by a pronounced Drude response (a peak centered at zero frequency) which is a hallmark of metallic behavior, no Drude response is observed in $\sigma_{1}^{c}(\omega)$. Instead, the far-infrared $\sigma_{1}^{c}(\omega)$ is dominated by sharp phonon peaks, indicating a semiconducting nature of La4Ni3O10 along the $c$ axis. Such semiconducting behavior in $\sigma_{1}^{c}(\omega)$ may be ascribed to the existence of insulating rock salt LaO layers, which effectively prevent the formation of coherence between different trilayers along the $c$ directions. The striking difference between $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ points to highly anisotropic charge dynamics in La4Ni3O10. The zero-frequency extrapolation of $\sigma_{1}(\omega)$ yields the dc conductivity $\sigma_{1}$($\omega\rightarrow$ 0) $\equiv 1/\rho_{\textrm{dc}}$. The inset of Fig. 1(b) plots $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K on a logarithmic scale; the red and blue solid circles on the vertical axis denote $\sigma_{1}^{ab}$($\omega\rightarrow$0) $\simeq$ 5490 $\Omega^{-1}\textrm{cm}^{-1}$ and $\sigma_{1}^{c}$($\omega\rightarrow$0) $\simeq$ 15 $\Omega^{-1}\textrm{cm}^{-1}$, respectively, which result in a resistivity anisotropy of $\rho_{c}$$\rho_{ab}$ $\simeq$ 366 at 300 K. This value is significantly larger than those of iron-based superconductors (FeSCs), such as FeSe ($\sim$3–4) Vedeneev et al. (2013), LiFeAs ($\sim$1–3) Song et al. (2010), and the BaFe2As2 family ($\sim$2–6) Tanatar et al. (2009a, b); Nakajima et al. (2018), but comparable to those of high-$T_{c}$ cuprates Komiya et al. (2002); Watanabe et al. (1997). Figure 1(d) compares $\sigma^{ab}_{1}(\omega)$ (red solid curve) and $\sigma^{c}_{1}(\omega)$ (blue solid curve) over a broad frequency range at 300 K, revealing that the interband electronic transitions are also strongly anisotropic in La4Ni3O10.
We fit the experimental $\sigma_{1}(\omega)$ to the Drude-Lorentz model (see Supplemental Materials Sup and Refs. Dressel and Grüner (2002); Tanner (2019)). Figure 2(a) displays the measured $\sigma_{1}^{ab}(\omega)$ at 150 K (cyan solid curve) and the fitting result (black dashed line), which is decomposed into two Drude components (D1 and D2) associated with intraband transitions, and a series of Lorentzian oscillators (L1ab, L2ab, L3ab and LHab) originating from interband transitions. In contrast, the fit of $\sigma_{1}^{c}(\omega)$ at 150 K [Fig. 2(b)] only requires several Lorentz components (L1c, L2c and LHc) to describe the interband transitions, as Drude behavior is absent in the $\sigma_{1}^{c}(\omega)$ spectrum of La4Ni3O10.
In order to elucidate the origin of these components in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$, we calculated the orbital-resolved electronic band structure of La4Ni3O10 (Supplemental Materials Sup ). As shown in Fig. 2(e), the bands near $E_{\textrm{F}}$ (within $\pm$2 eV) are dominated by Ni-$d_{x^{2}-y^{2}}$ (navy) and Ni-$d_{z^{2}}$ (magenta) orbitals, similar to the bilayer La3Ni2O7 Abadi et al. (2025); Christiansson et al. (2023); Geisler et al. (2024); LaBollita et al. ; Liu et al. (2024); Li et al. (2024c); Luo et al. (2023); Sun et al. (2023b); Yang et al. (2024a); Zhang et al. (2023). Due to strong hybridization between the Ni-$d_{z^{2}}$ and O-$p_{z}$ orbitals, the bands originating from the Ni-$d_{z^{2}}$ orbital are split into anti-bonding (AB), non-bonding (NB) and bonding (B) branches Li et al. (2024b); Sakakibara et al. (2024); Chen et al. (2024c); Zhang et al. (2024c, d), whereas the Ni-$d_{x^{2}-y^{2}}$ orbitals lying in the NiO6 plane are only weakly coupled between planes, thus not undergoing the AB-NB-B splitting. While the multiple bands formed by the Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals crossing $E_{\textrm{F}}$ give rise to the two Drude components in $\sigma_{1}^{ab}(\omega)$, electronic transitions between these bands are responsible for the Lorentz components in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$. To gain further insights into the interband transitions, we calculated the energy- and momentum-resolved transition dipole moments (TDMs) $P_{\alpha}^{2}(n,\vec{k})$ for $E\parallel ab$ [Fig. 2(f)] and $E\parallel c$ [Fig. 2(g)] (Supplemental Materials Sup and Refs. Geisler et al. (2024)). A combination of the TDMs and the orbital-resolved band structure [Fig. 2(e)] allows us to clarify the origin of the Lorentz components in $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$. For $E\parallel ab$, the interband transitions mainly occur between the Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals: L1ab stems from the transitions between the occupied Ni-$d_{z^{2}}$ bonding states and the empty Ni-$d_{x^{2}-y^{2}}$ bands (green arrow); L2ab arises from transitions from the occupied Ni-$d_{x^{2}-y^{2}}$ bands to the empty Ni-$d_{z^{2}}$ non-bonding bands and from the occupied Ni-$d_{z^{2}}$ bonding bands to the empty Ni-$d_{x^{2}-y^{2}}$ bands (orange arrows); the transitions from the occupied Ni-$d_{x^{2}-y^{2}}$ bands to the empty Ni-$d_{z^{2}}$ anti-bonding bands and from the occupied Ni-$d_{z^{2}}$ bonding bands to the empty Ni-$d_{x^{2}-y^{2}}$ bands (violet arrows) give rise to L3ab. For $E\parallel c$, the interband excitations are dominated by electronic transitions between Ni-$d_{z^{2}}$ orbitals: both L1c and L2c originate from electronic transitions between the occupied Ni-$d_{z^{2}}$ bonding and the empty Ni-$d_{z^{2}}$ non-bonding bands, as denoted by the green and orange arrows in Fig. 2(g). Both LHab and LHc correspond to high-energy transitions involving bands far from $E_{\textrm{F}}$.
Figures 2(c) and 2(d) show the calculated interband $\sigma_{1}(\omega)$ for $E\parallel ab$ and $E\parallel c$, respectively Sup . While the calculated $\sigma_{1}(\omega)$ qualitatively captures the interband components in the measured $\sigma_{1}(\omega)$ for both polarizations, there are noticeable discrepancies between the calculated and experimental $\sigma_{1}(\omega)$. Such discrepancies arise from the effects of electronic correlations that are not taken into account in the DFT calculations. In correlated materials, electronic correlations lead to a renormalization of the electronic bands, i.e. a reduction of the band width, which induces a shift of the interband transition peaks in the experimental $\sigma_{1}(\omega)$ towards lower energy compared to DFT calculations Qazilbash et al. (2009); Si (2009); Xu et al. (2020); Cao et al. (2025); Yi et al. (2025). Here in La4Ni3O10, the Lorentz components in the measured $\sigma_{1}^{ab}(\omega)$ [Fig. 2(a)] and $\sigma_{1}^{c}(\omega)$ [Fig. 2(b)] all exhibit a noticeable redshift compared to the interband transition peaks in the calculated $\sigma_{1}^{ab}(\omega)$ [Fig. 2(c)] and $\sigma_{1}^{c}(\omega)$ [Fig. 2(d)], indicating that the bands formed by Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals are renormalized due to electronic correlation effects. It is also noteworthy that the redshift for each interband component is different from others, implying that the electronic-correlation-induced renormalization varies from band to band in La4Ni3O10. Furthermore, L3ab and L1c in the measured $\sigma_{1}(\omega)$ are significantly smaller or broader than the corresponding interband transition peaks in the calculated $\sigma_{1}(\omega)$, whereas other components agree relatively better with theoretical calculations. This fact also points to band-dependent electronic correlation effects. ARPES measurements Li et al. (2017); Du et al. and DFT+DMFT calculations Wang et al. (2024b) have found that the Ni-$d_{z^{2}}$ orbital features stronger electronic correlations than the Ni-$d_{x^{2}-y^{2}}$ orbital, compatible with our optical results.
Figure 3: The $S$ ratios between different temperatures as a function of $\omega_{c}$ for (a) $E\parallel ab$ and (b) $E\parallel c$. (c)-(h) The $T$ dependence of $S$ (normalized by its value at 300 K) for different cutoff frequencies.
Next, we investigate the $T$ dependence of the charge dynamics by tracking the transfer of the spectral weight defined as
$S=\int_{0}^{\omega_{c}}\sigma_{1}(\omega)d\omega,$ (1)
where $\omega_{c}$ is a cutoff frequency. Figures 3(a) and 3(b) display the $S$ ratios between different temperatures for $E\parallel ab$ and $E\parallel c$, respectively, which conveniently reflect the direction and energy scale of spectral weight transfer. Specifically, an $S$ ratio exceeding 1 corresponds to a spectral weight transfer from high to low energy, whereas a spectral weight transfer from low to high energy leads to an $S$ ratio lower than 1. When $\omega_{c}$ covers the full energy scale of the spectral weight transfer, the $S$ ratio reaches 1 due to the conservation of spectral weight.
$S^{ab}(150\penalty 10000\ \text{K})/S^{ab}(300\penalty 10000\ \text{K})$ is larger than 1 and approaches 1 with increasing $\omega_{c}$. Although $S^{c}(150\penalty 10000\ \text{K})/S^{c}(300\penalty 10000\ \text{K})$ is below 1 in the low-frequency range, which is most likely associated with the reduction of thermally excited carriers with decreasing $T$, it exceeds 1 for $\omega_{c}>2000$ $\textrm{cm}^{-1}$. These observations suggest that in the pristine phase, as $T$ is reduced, there is a spectral weight transfer from high to low energy over a broad frequency range in both $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$. This effect can be further confirmed by the $T$ dependence of $S^{ab}$ and $S^{c}$: as shown in Figs. 3(c)-3(h), for different $\omega_{c}$’s, both $S^{ab}(T)/S^{ab}(300\penalty 10000\ \text{K})$ and $S^{c}(T)/S^{c}(300\penalty 10000\ \text{K})$ increase monotonically as $T$ decreases from 300 K to 150 K. In a Hund’s metal, the decrease of $T$ results in a spectral weight transfer from low to high frequency. Such behavior has been reported in iron-based superconductors and attributed to the strong Hund’s coupling between itinerant electrons and localized electron moments Schafgans et al. (2012); Wang et al. (2012); Wu et al. (2011); Xu et al. (2014). On the contrary, Mott-Hubbard systems are expected to feature a spectral weight transfer from high to low frequency as $T$ is lowered Kotliar and Vollhardt (2004); Toschi et al. (2005), which has been widely observed in cuprates Molegraaf et al. (2002); Toschi et al. (2005); Ortolani et al. (2005); Carbone et al. (2006a, b); Santander-Syro et al. (2002), V2O3 in the high-temperature metallic state Qazilbash et al. (2008), and Nd4Ni3O8 Hao et al. (2021). The spectral weight redistribution in La4Ni3O10 hints that Coulomb correlations exert considerable influence on the charge dynamics which place La4Ni3O10 closer to a Mott system rather than a Hund’s system. It is also worth noting that in the pristine phase, the $T$ dependence of $S^{ab}$ and $S^{c}$ for all different $\omega_{c}$’s follows $T^{2}$ behavior [blue solid lines in Figs. 3(c)-3(h)], in accord with the calculations based on the dynamical mean-field theory of the Hubbard model Toschi et al. (2005) and the experimental results in cuprates Molegraaf et al. (2002); Ortolani et al. (2005); Carbone et al. (2006a, a); Hao et al. (2021) and Nd4Ni3O8 Hao et al. (2021).
While $S^{ab}(5\penalty 10000\ \text{K})/S^{ab}(150\penalty 10000\ \text{K})$ exceeds 1 in the very low-frequency range, which is due to the narrowing of the Drude response, it falls below 1 for $\omega_{c}>500$ $\textrm{cm}^{-1}$; $S^{c}(5\penalty 10000\ \text{K})/S^{c}(150\penalty 10000\ \text{K})$ lies below 1 and approaches 1 as $\omega_{c}$ grows. These observations unambiguously suggest that for both $E\parallel ab$ and $E\parallel c$, a spectral weight transfer from low to high energy occurs at low temperatures. Furthermore, Figs. 3(c)-3(h) show that for different $\omega_{c}$’s, both $S^{ab}(T)/S^{ab}(300\penalty 10000\ \text{K})$ and $S^{c}(T)/S^{c}(300\penalty 10000\ \text{K})$ exhibit a drop below the density-wave transition at $T_{\text{DW}}$, attesting to the suppression of low-frequency spectral weight. All these results are in agreement with the optical response of a density-wave gap, which suppresses the low-frequency spectral weight, i.e. the Drude weight in $\sigma_{1}(\omega)$, and transfers it to higher frequency Hu et al. (2008); Zhou et al. (2021); Liu et al. (2024); Xu et al. (2025). Here, it is worth noting that for $E\parallel c$, Drude behavior is absent, suggesting that the spectral weight transfer in $\sigma_{1}^{c}(\omega)$ originates from a redistribution of interband transitions. Since $\sigma_{1}^{c}(\omega)$ is dominated by interband transitions between the occupied Ni-$d_{z^{2}}$ bonding states and the empty Ni-$d_{z^{2}}$ non-bonding states [Fig. 2(g)], the observation of density-wave-induced spectral weight redistribution in $\sigma_{1}^{c}(\omega)$ implies that the Ni-$d_{z^{2}}$ orbital is involved in the formation of the density-wave gap, highlighting the important role of the Ni-$d_{z^{2}}$ orbital in driving the density-wave instabilities. Note that previous ARPES studies on La4Ni3O10 revealed contradictory results regarding the question of which bands are gapped by the density-wave transition Li et al. (2017); Du et al. , calling for further investigations into the momentum dependence of the density-wave gap.
In summary, the abab-plane and $c$-axis charge dynamics of La4Ni3O10 have been studied using optical spectroscopy. While $\sigma_{1}^{ab}(\omega)$ exhibits a pronounced Drude peak, indicating a metallic nature of La4Ni3O10 in the abab plane, semiconducting behavior is observed in $\sigma_{1}^{c}(\omega)$. The zero-frequency extrapolation of $\sigma_{1}^{ab}(\omega)$ and $\sigma_{1}^{c}(\omega)$ at 300 K yields a resistivity anisotropy of $\rho_{c}/\rho_{ab}\simeq 366$ for La4Ni3O10, which is much higher than the values in FeSCs, but similar to those in cuprates. The interband response also features strong anisotropy, as light polarized in the abab plane and along the $c$ axis selectively excites electronic transitions between different orbitals. In the pristine phase, the spectral weight transfer implies that the charge dynamics is under the influence of Coulomb correlations, whereas in the density-wave state, the spectral weight redistribution suggests the opening of a gap with the participation of the Ni-$d_{z^{2}}$ orbital.
Acknowledgements.
We thank Ilya M. Eremin, Frank Lechermann, Qianghua Wang and Shunli Yu for helpful discussions. Work at NJU was supported by the National Key R&D Program of China (Grants No. 2022YFA1403201 and 2022YFA1403000), the National Natural Science Foundation of China (Grants No. 12174180 and 12574144). Work at SYSU was supported by the National Natural Science Foundation of China (Grant No. 12425404), the National Key Research and Development Program of China (Grant No. 2023YFA1406500), the CAS Superconducting Research Project (Grant No. SCZX-0101), the Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (Grant No. 2022B1212010008), and Research Center for Magnetoelectric Physics of Guangdong Province (Grant No. 2024B0303390001).